[Hallc_running] [New Logentry] Target pressure issue fixed, plan to take data on LH2

alis at jlab.org alis at jlab.org
Sat Dec 11 14:20:02 EST 2021

Logentry Text:
Earlier in the day shift it was noticed that the pressure of the LH2 target was very low, 16-18 psi which is unacceptable. Around 1:30 pm-2:00 pm, we went into controlled access, and Dave Meekins went in the hall to investigate, and noticed one of the main supply valves (connected to ballast tanks) was off, and he had to manually turn it on. The pressure is now stable for LH2, and we are ready to take data for LH2. The pressure limits for the alarm handler have also been adjusted to 20-22 psi (previously they were 14-17 psi..).


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3960934
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20211211/bea89cb0/attachment.html>

More information about the Hallc_running mailing list